How many diagonals can be drawn in the interior of a regular polygon if one of its exterior angles contain 60 degrees?
Answer by stanbon(75887)

How many diagonals can be drawn in the interior of a regular polygon if one of its exterior angles contain 60 degrees?
-------
Sum of exterior angles = 360 degrees.
# of exteror angles = # of sides = 360/60 = 6
---------
# of sides + # of diagonals:: 6C2 = 15
----
# of diagonals = 15-6 = 9
==========
Cheers,
Stan H.